OpenStudy (anonymous):

What are the three measurements you need to make in order to calculate power? What are the units of those measurements?

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

I don't even know what it means by "power", what kind of power?

OpenStudy (gorv):

wher u got power ???

OpenStudy (gorv):

u heard abt work in physics ??

OpenStudy (anonymous):

It says in order to calculate, "power". I know the formula is Power=Work/Time

OpenStudy (anonymous):

Oh are those the measurements? x3

OpenStudy (gorv):

we need to find in term of M ,L ,T ??

OpenStudy (gorv):

yeah

OpenStudy (anonymous):

But that's solely 2, Work and Time.

OpenStudy (gorv):

work = force * distance

OpenStudy (anonymous):

That would be the unit it's measure in, right?

OpenStudy (gorv):

what ..i didnt get u ?? :P

OpenStudy (gorv):

(force*distance )/time

OpenStudy (anonymous):

I mean the measurements are Time and Work, but it's also asking for the units of those measurements.

OpenStudy (gorv):

yeah yeah we r coming to that dear

OpenStudy (anonymous):

So the measurements are Time, distance&force, and Work, I'm perplexed.

OpenStudy (gorv):

work /time = (kg*m^2/s^2)/sec

OpenStudy (gorv):

no force distance and time

OpenStudy (gorv):

=(kg*m^2)/s^3

OpenStudy (anonymous):

Ohhhh, I get it Work = distance * force

OpenStudy (gorv):

yeah yeah your right ....smart girl

OpenStudy (gorv):

can i go now ?/

OpenStudy (anonymous):

You haven't told me the units of the measurements.

OpenStudy (gorv):

=(kg*m^2)/s^3

OpenStudy (gorv):

or N*m/s

OpenStudy (gorv):

u can use both

OpenStudy (anonymous):

Thank you Gorv. <3

OpenStudy (gorv):

wlcm :)